# Strong Quantum Mpemba Effect with Squeezed Thermal Reservoirs

### v1

@@ Acknowledgements
JF would like to thank the Fundação Cearense de Apoio ao Desenvolvimento Científico e Tecnológico (FUNCAP) under the grant PRONEM PNE0112-00085.01.00/16 for financial support, and the Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq) under the grant 304485/2023-3. ACS acknowledges the support by the European Union’s Horizon 2020 FET-Open project SuperQuLAN (899354), and by the Proyecto Sinérgico CAM 2020 Y2020/TCS-6545 (NanoQuCo-CM) from the Comunidad de Madrid.
@@ A.1 The superoperator formalism
Our discussion is mainly derived from Refs. [28, 16, 29, 39]. First of all, we adopt a traceless matrix basis $\{\sigma_{1},\sigma_{2},\cdots,\sigma_{D^{2}-1}\}$ that satisfies $\mathrm{tr}[\sigma_{n}\sigma^{\dagger}_{m}]=D\delta_{nm}$, the density matrix can be written as
@@
Without loss of generality, here we consider the class of evolution where the system dynamics is described by the time-local master equation [30, 31]
@@ A.2 The superoperator spectrum
As an important remark, this change of formalism has a price to be paid, and it is mainly due to the non-Hermiticity of the superoperator $\mathbbm{L}$. Nevertheless, non-Hermitian operators can always be written in the Jordan canonical form, where $\mathbbm{L}(t)$ displays a block-diagonal structure $\mathbbm{L}_{\mathrm{J}}(t)$ composed by Jordan blocks $J_{n}(t)$ associated with different time-dependent eigenvalues $\lambda_{\alpha}(t)$ of $\mathbbm{L}(t)$ [40]. When such a Jordan decompoistion is possible, we can write the diagonal form of the superoperator $\mathbbm{L}$ as
@@ Appendix B Superoperator formalism and Mpemba effect
because $\mathrm{Re}[\lambda_{n}]<0$, and therefore we can observe the emergence of the Mpemba effect only considering the time-dependent term in equation above. In this way, we stablish some sufficient conditions to observe strong Mpemba effect. By ordering $\mathrm{Re}[\lambda_{1}]\leq\mathrm{Re}[\lambda_{2}]\leq\mathrm{Re}[\lambda_{3}]$, we have the conditions to be satisfied for the hot initial state $|\psi_{\mathrm{hot}}(0)\rangle\rangle$

### v2

@@ Acknowledgements
JF would like to thank the Fundação Cearense de Apoio ao Desenvolvimento Científico e Tecnológico (FUNCAP) under the grant PRONEM PNE0112-00085.01.00/16 for financial support, and the Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq) under the grant 304485/2023-3. ACS is supported by the Comunidad de Madrid through the program Talento 2024 “César Nombela”, under Grant No 2024-T1/COM-31530 (project SWiQL). ACS also acknowledges the support from the European Union’s Horizon 2020 FET-Open project SuperQuLAN (899354) and the Proyecto Sinérgico CAM 2020 Y2020/TCS-6545 (NanoQuCo-CM) from the Comunidad de Madrid.
@@ A.1 The superoperator formalism
Our discussion is mainly derived from Refs. Sarandy:05-1 ; Sarandy:05-2 ; Santos:20d ; Santos:21c . First of all, we adopt a traceless matrix basis $\{\sigma_{1},\sigma_{2},\cdots,\sigma_{D^{2}-1}\}$ that satisfies $\mathrm{tr}[\sigma_{n}\sigma^{\dagger}_{m}]=D\delta_{nm}$, the density matrix can be written as
@@
Without loss of generality, here we consider the class of evolution where the system dynamics is described by the time-local master equation Alicki:Book07 ; Petruccione:Book
@@ A.2 The superoperator spectrum
As an important remark, this change of formalism has a price to be paid, and it is mainly due to the non-Hermiticity of the superoperator $\mathbbm{L}$. Nevertheless, non-Hermitian operators can always be written in the Jordan canonical form, where $\mathbbm{L}(t)$ displays a block-diagonal structure $\mathbbm{L}_{\mathrm{J}}(t)$ composed by Jordan blocks $J_{n}(t)$ associated with different time-dependent eigenvalues $\lambda_{\alpha}(t)$ of $\mathbbm{L}(t)$ Horn:Book . When such a Jordan decompoistion is possible, we can write the diagonal form of the superoperator $\mathbbm{L}$ as
@@ Appendix B Superoperator formalism and Mpemba effect
because $\mathrm{Re}[\lambda_{n}]<0$, and therefore we can observe the emergence of the Mpemba effect only considering the time-dependent term in equation above. In this way, we stablish some sufficient conditions to observe different regimes of Mpemba effect. By ordering $\mathrm{Re}[\lambda_{1}]\leq\mathrm{Re}[\lambda_{2}]\leq\mathrm{Re}[\lambda_{3}]$, we have the conditions to be satisfied for the hot initial state $|\psi_{\mathrm{hot}}(0)\rangle\rangle$